OpenStudy (anonymous):

What is 0.47% written as a decimal?

OpenStudy (anonymous):

take away the percent

OpenStudy (anonymous):

0.47 :)

OpenStudy (anonymous):

.0047

OpenStudy (anonymous):

0.47 for sure

OpenStudy (anonymous):

you have to move the number back to convert it back to a decimal so it would be .0047

OpenStudy (anonymous):

0.47 %=0.47/100 \[\frac{ 0.47 }{ 100 }=\frac{ 47 }{ 100\times100 }=0.0047\]

OpenStudy (anonymous):

Actually I think it is 0.0047. Yeah because 0.0047 as a percent is 47%.